# Language-Guided Image Tokenization for Generation

### v1

@@ Abstract
Image tokenization, the process of transforming raw image pixels into a compact low-dimensional latent representation, has proven crucial for scalable and efficient image generation. However, mainstream image tokenization methods generally have limited compression rates, making high-resolution image generation computationally expensive. To address this challenge, we propose to leverage language for efficient image tokenization, and we call our method Text-Conditioned Image Tokenization (TexTok). TexTok is a simple yet effective tokenization framework that leverages language to provide high-level semantics. By conditioning the tokenization process on descriptive text captions, TexTok allows the tokenization process to focus on encoding fine-grained visual details into latent tokens, leading to enhanced reconstruction quality and higher compression rates. Compared to the conventional tokenizer without text conditioning, TexTok achieves average reconstruction FID improvements of 29.2% and 48.1% on ImageNet-256 and -512 benchmarks respectively, across varying numbers of tokens. These tokenization improvements consistently translate to 16.3% and 34.3% average improvements in generation FID. By simply replacing the tokenizer in Diffusion Transformer (DiT) with TexTok, our system can achieve a 93.5$\times$ inference speedup while still outperforming the original DiT using only 32 tokens on ImageNet-512. TexTok with a vanilla DiT generator achieves state-of-the-art FID scores of 1.46 and 1.62 on ImageNet-256 and -512 respectively. Furthermore, we demonstrate TexTok’s superiority on the text-to-image generation task, effectively utilizing the off-the-shelf text captions in tokenization.
@@ 1 Introduction
Addressing this limitation requires a fundamental shift in the tokenization process. At its core, tokenization involves finding a compact and effective representation of an image. The most concise and meaningful representation of an image often comes from its language description–i.e., captioning. When describing an image, humans naturally start with high-level semantics before elaborating on finer details. Inspired by this insight, we introduce Text-Conditioned Image Tokenization (TexTok), a novel framework that leverages descriptive text captions to provide high-level semantic content, and thus the tokenization process focuses more on fine-grained visual details, thereby enhancing reconstruction quality without compromising compression rate.
@@ 2 Related Work
Many recent studies start to leverage external semantic information, such as image representations and semantic maps, to improve image generation [33, 27, 51]. Unlike these methods, which use external semantics to aid the generation process, our approach focuses on enhancing the tokenization process through conditioning on text semantics. Some recent efforts [48, 30, 53, 29] also consider aligning image tokens with text semantics in image tokenization to improve multimodal understanding. They either directly map images to text tokens in a frozen LLM codebook [48, 30, 53] or align the features of image tokens with text features [29], to produce semantically meaningful tokens. However, by enforcing strict image-text alignments, these works suffer from limited image reconstruction quality due to the inherent divergence between vision and language representations, resulting in undesirable image generation quality. In contrast, our work takes a complementary approach. We leverage text captions as external semantic conditioning, allowing image tokens to focus on capturing finer visual details, significantly boosting the image reconstruction and generation performance.
@@ 3.2 TexTok: Text-Conditioned Image Tokenization
We introduce Text-Conditioned Image Tokenization (TexTok), a simple yet effective tokenization framework that leverages language to provide high-level semantics and focuses more on tokenizing fine-grained visual details. Unlike existing methods that compress all visual information into latent tokens, we use descriptive text captions to represent high-level semantics and guide the tokenization process.
@@
Given an image caption, we use a frozen T5 [35] text encoder to extract text embeddings. These embeddings are injected into both the tokenizer and detokenizer, providing semantic guidance throughout the tokenization process and allowing the learned tokens to focus more on capturing fine-grained visual details.
@@
By directly injecting text tokens containing high-level semantic information into both the tokenizer and detokenizer, TexTok alleviates the need for the tokenizer and detokenizer to learn the semantics, enabling them to focus more on encoding the remaining fine-grained visual details into the image tokens, which greatly improves the compression rate without sacrificing reconstruction performance.
@@
During inference, the process differs by the generation task. For text-to-image generation, we use the provided captions for both tokenization and generation, feeding the text embeddings and generated latent image tokens into the detokenizer to produce the output image. For class-conditional generation, DiT generates latent tokens based on the specified class; we then sample an unseen caption for that class from a pre-generated list, and inject it into the detokenizer along with the generated latent tokens to produce the final image. Notably, only the class category is used during generation, in line with standard practice.
@@ 4.1 Implementation Details
We use Diffusion Transformer (DiT) [32] as our default generator due to its effectiveness and flexibility of handling 1D tokens. We use a DiT patch size of 1 for all TexTok generation experiments, and by default, we train DiT for 350 epochs. Specifically, for class-conditional generation, we use the original DiT architecture. For text-to-image generation, referring to [5], we modify DiT architecture by adding an additional multi-head cross-attention layer following the multi-head self-attention layer in the DiT block to accept text embeddings. We refer to this architecture as “DiT-T2I”.
@@ 4.2 Experiment Setup
We compare two setups to demonstrate the effectiveness of using text conditioning: TexTok incorporates text tokens in both the tokenizer and detokenizer, corresponding to the architecture shown in Figure 2. In contrast, Baseline (w/o text) does not condition on text tokens in both the tokenizer and detokenizer. For each image, we tokenize it into “#tokens” number of latent tokens and train the generator to generate these tokens.
@@ 4.4 System-level Image Generation Comparison
Table 2: System-level comparison of class-conditional image generation on ImageNet 256$\times$256 and 512$\times$512. TexTok-256 + DiT-XL achieves state-of-the-art performance on both image resolutions. All entries use classifier-free guidance if applicable. Note that our method is orthogonal to both latent generation models and classifier-free guidance techniques, more advanced latent generators [31] and guidance mechanisms [25, 20] can also be applied to TexTok to further improve our performance. “#Params (G)”: the number of generator’s parameters. “#Params (T)”: the number of tokenizer’s parameters. “#tokens”: the number of latent image tokens used during generation. “/” in the first three columns indicates different configurations used at different image resolutions respectively. denotes the number of tokens before patchification.
@@ 5 Conclusion
We present Text-Conditioned Image Tokenization (TexTok), a framework that leverages descriptive text captions to provide high-level semantics, allowing tokenization to focus on encoding fine-grained visual details into latent tokens. TexTok significantly improves reconstruction and generation performance, achieving state-of-the-art results on ImageNet with high computational efficiency. By mitigating the trade-off between reconstruction quality and compression rates, TexTok enables more efficient image generation.
@@ Acknowledgements.
We thank David Minnen, Eirikur Agustsson, Qihang Yu, Peng Cao, José Lezama, Long Zhao, Haotian Tang, Tianhong Li, Xuhui Jia, Ruben Villegas and Xingyi Zhou for helpful discussions and valuable feedback.
@@ Appendix F Additional Discussion on Related Work
In contrast, our work introduces a novel image tokenization framework specifically designed for generative tasks. We are the first work that proposes conditioning the tokenization process directly on the text captions of images, a strategy typically reserved for the generation phase. Rather than enforcing a strict alignment between image tokens and text captions, our method leverages descriptive text captions to provide high-level semantics while allowing image tokens to focus on capturing fine-grained visual details. This complementary approach significantly improves both reconstruction quality and compression rate of tokenization. Furthermore, we demonstrate that our approach achieves state-of-the-art performance on standard ImageNet conditional generation benchmarks on both 256$\times$256 and 512$\times$512 image resolutions, establishing our method as a distinct advancement over existing works.

### v2

@@ Abstract
Image tokenization, the process of transforming raw image pixels into a compact low-dimensional latent representation, has proven crucial for scalable and efficient image generation. However, mainstream image tokenization methods generally have limited compression rates, making high-resolution image generation computationally expensive. To address this challenge, we propose to leverage language for efficient image tokenization, and we call our method Text-Conditioned Image Tokenization (TexTok). TexTok is a simple yet effective tokenization framework that leverages language to provide a compact, high-level semantic representation. By conditioning the tokenization process on descriptive text captions, TexTok simplifies semantic learning, allowing more learning capacity and token space to be allocated to capture fine-grained visual details, leading to enhanced reconstruction quality and higher compression rates. Compared to the conventional tokenizer without text conditioning, TexTok achieves average reconstruction FID improvements of 29.2% and 48.1% on ImageNet-256 and -512 benchmarks respectively, across varying numbers of tokens. These tokenization improvements consistently translate to 16.3% and 34.3% average improvements in generation FID. By simply replacing the tokenizer in Diffusion Transformer (DiT) with TexTok, our system can achieve a 93.5$\times$ inference speedup while still outperforming the original DiT using only 32 tokens on ImageNet-512. TexTok with a vanilla DiT generator achieves state-of-the-art FID scores of 1.46 and 1.62 on ImageNet-256 and -512 respectively. Furthermore, we demonstrate TexTok’s superiority on the text-to-image generation task, effectively utilizing the off-the-shelf text captions in tokenization. Project page is at: https://kaiwenzha.github.io/textok/.
@@ 1 Introduction
Addressing this limitation requires a fundamental shift in the tokenization process. At its core, tokenization involves finding a compact and effective representation of an image. The most concise and meaningful representation of an image often comes from its language description–i.e., captioning. When describing an image, humans naturally start with high-level semantics before elaborating on finer details. Inspired by this insight, we introduce Text-Conditioned Image Tokenization (TexTok), a novel framework that leverages text captions to guide the tokenizer in learning image semantics. This simplifies semantic learning, allowing more learning capacity and token space to be allocated to capture fine-grained visual details, thereby enhancing reconstruction quality without compromising compression rate.
@@ 2 Related Work
Many recent studies start to leverage external semantic information, such as image representations and semantic maps, to improve image generation [33, 27, 51]. Unlike these methods, which use external semantics to aid the generation process, our approach focuses on enhancing the tokenization process through conditioning on text semantics. Some recent efforts [48, 30, 53, 29] also consider aligning image tokens with text semantics in image tokenization to improve multimodal understanding. They either directly map images to text tokens in a frozen LLM codebook [48, 30, 53] or align the features of image tokens with text features [29], to produce semantically meaningful tokens. However, by enforcing strict image-text alignments, these works suffer from limited image reconstruction quality due to the inherent divergence between vision and language representations, resulting in undesirable image generation quality. In contrast, our work takes a complementary approach. We leverage text as external semantic conditioning, significantly boosting the image reconstruction and generation performance.
@@ 3.2 TexTok: Text-Conditioned Image Tokenization
We introduce Text-Conditioned Image Tokenization (TexTok), a simple yet effective tokenization framework. Unlike existing methods that compress all visual information into latent tokens, we use text captions to represent high-level semantics and guide the tokenization process.
@@
Given an image caption, we use a frozen T5 [35] text encoder to extract text embeddings. These embeddings are injected into both the tokenizer and detokenizer to providing semantic guidance throughout the tokenization process.
@@
By directly injecting text tokens containing high-level semantic information into both the tokenizer and detokenizer, TexTok alleviates the need for the tokenizer and detokenizer to learn image semantics.
@@
During inference, the process differs by the generation task. For text-to-image generation, we use the provided captions for both generation and detokenization, feeding the text embeddings and generated latent image tokens into the detokenizer to produce the output image. For class-conditional generation, DiT generates latent tokens based on the specified class; we then sample an unseen caption for that class from a pre-generated list, and inject it into the detokenizer along with the generated latent tokens to produce the final image. Notably, only the class category is used during generation, in line with standard practice.
@@ 4.1 Implementation Details
We use Diffusion Transformer (DiT) [32] as our default generator due to its effectiveness and flexibility of handling 1D tokens. We use a DiT patch size of 1 for all TexTok generation experiments, and by default, we train DiT for 350 epochs. Specifically, for class-conditional generation, we use the original DiT architecture. For text-to-image generation, referring to [5], we modify DiT architecture by adding an additional multi-head cross-attention layer following the multi-head self-attention layer in the DiT block to accept text embeddings. We refer to this architecture as “DiT-T2I".
@@ 4.2 Experiment Setup
We compare two setups to demonstrate the effectiveness of using text conditioning: TexTok incorporates text tokens in both the tokenizer and detokenizer, corresponding to the architecture shown in Figure 2. In contrast, Baseline (w/o text) does not condition on text tokens in both the tokenizer and detokenizer. For each image, we tokenize it into “#tokens" number of latent tokens and train the generator to generate these tokens.
@@ 4.4 System-level Image Generation Comparison
Table 2: System-level comparison of class-conditional image generation on ImageNet 256$\times$256 and 512$\times$512. TexTok-256 + DiT-XL achieves state-of-the-art performance on both image resolutions. All entries use classifier-free guidance if applicable. Note that our method is orthogonal to both latent generation models and classifier-free guidance techniques, more advanced latent generators [31] and guidance mechanisms [25, 20] can also be applied to TexTok to further improve our performance. “#Params (G)": the number of generator’s parameters. “#Params (T)": the number of tokenizer’s parameters. “#tokens": the number of latent image tokens used during generation. “/" in the first three columns indicates different configurations used at different image resolutions respectively. denotes the number of tokens before patchification.
@@ 5 Conclusion
We present Text-Conditioned Image Tokenization (TexTok), a new framework that leverages captions to guide the tokenizer in learning image semantics, allowing more learning capacity and token space to be allocated to capture visual details. TexTok significantly improves both reconstruction and generation performance, achieving state-of-the-art results in conditional and text-to-image generation on ImageNet with computational efficiency. By mitigating the trade-off between reconstruction quality and compression rate, TexTok enables more efficient image generation.
@@ Acknowledgements.
We thank David Minnen, Eirikur Agustsson, Qihang Yu, Peng Cao, José Lezama, Long Zhao, Haotian Tang, Tianhong Li, Xuhui Jia, Ruben Villegas and Xingyi Zhou for helpful discussions and valuable feedback. KZ and DK are partly funded by Wistron Corporation.
@@ Appendix F Additional Discussion on Related Work
In contrast, our work introduces a novel image tokenization framework specifically designed for generative tasks. We are the first work that proposes conditioning the tokenization process directly on the text captions of images, a strategy typically reserved for the generation phase. Rather than enforcing a strict alignment between image tokens and text captions, our method leverages descriptive captions to provide a compact semantic representation. This simplifies semantic learning, allowing more learning capacity and token space to be allocated to capture fine-grained visual details. This complementary approach significantly improves the reconstruction quality and compression rate of tokenization. Furthermore, we demonstrate that our approach achieves state-of-the-art performance on standard ImageNet conditional generation benchmarks on both 256$\times$256 and 512$\times$512 image resolutions, establishing our method as a distinct advancement over existing works.